Roland Stigge Debian Developer address placeholder stigge@debian.org

Roland Stigge, Debian Project

University of Leeds Computer Based Learning Unit LEEDS, LS2 9JT United Kingdom

9. Januar 2004

Subject: The license of L^{A}TEX2HTML

Dear Sir or Madam,

I’m a developer of the Free Software / Open Source project called Debian^{1} and personally maintain the popular package L^{A}TEX2HTML^{2} for it. This conversion tool, designed to translate from the L^{A}TEX publishing format to HTML, was initially written in 1993 by Mr Nikos Drakos during his work with the Computer Based Learning Unit at the University of Leeds.

The reason I am writing to you is that there is some uncertainty about the ownership of the copyright that vests in L^{A}TEX2HTML. Mr Drakos has signalled a willingness to change the licensing terms for L^{A}TEX2HTML but wonders if he needs your permission or support to do so due to the fact that he was employed by the University at the time of creating the software. As a precaution, we would like your written agreement to change the license terms for L^{A}TEX2HTML from its current license to the GNU General Public License (GPL)^{3}.

The current license for L^{A}TEX2HTML was unquestionably written as a Free Software / Open Source license, however one aspect of the license causes the Debian project difficulties. The second clause reads:

o  No fees or compensation are charged for use, copies, or
   access to this software. You may charge a nominal
   distribution fee for the physical act of transferring a
   copy, but you may not charge for the program itself.

This violates our self-imposed “Debian Free Software Guidelines” (DFSG^{4}), because it doesn’t allow the software to be sold. Debian is actually one of the least commercial system distributions (not even directly providing physical CDs, DVDs or support contracts) and is a wholly volunteer project as e.g. Linux itself. However, a common way for users to acquire the Debian system is to purchase it on CD or DVD from an independent vendor.

1See http://www.debian.org/ 2http://www.latex2html.org/ 3Verbatim copies of these two licenses are enclosed.

4http://www.de.debian.org/social contract#guidelines

For this reason, the Debian Project has concluded that the current license does not comply with our Guidelines^{5} and L^{A}TEX2HTML has been removed from Debian’s main distribution. Changing the license terms would allow the program to return to our main distribution and facilitate the already existing large user base of L^{A}TEX2HTML.

There are a number of arguments in favour of changing the license:

• The original author of the package, Mr Drakos, has signalled some willingness to change the license, and only wonders if he needs the support of the University of Leeds, which was his employer at the time he created it. Ross Moore, the current maintainer of L^{A}TEX2HTML, also expressed willingness to cooperate.

• The GPL is a popular license in the area of Free Software / Open Source. In contrast to individual licenses like the current one for L^{A}TEX2HTML, it is well known around the world and under constant review by the user and developer base to prevent legal problems (possibly like the one in question).

• As you will see in the current license itself, some parts of L^{A}TEX2HTML are already licensed under the protection of the GPL. This probably means that the whole package should be licensed under the GPL anyway.

• Debian is not the only project distributing L^{A}TEX2HTML. Other distributions of the popular GNU/Linux system like RedHat and SuSE and different BSD derivates also use L^{A}TEX2HTML and probably have the same problem.

• The change would protect the investment of the University of Leeds in the Free Software community by ensuring further widespread use of L^{A}TEX2HTML.

If you have any queries, please do not hesitate to contact me. I look forward to your reply. Thanks in advance.

Sincerely,

Roland Stigge Debian Developer

5http://lists.debian.org/debian-legal/2003/debian-legal-200310/msg00383.html and follow-ups